OpenStudy (hrazor):

A piece of tin is in the form of a rectangle whose area is 486 sq. cm. A square side 3 cm is cut from each corner, and an open box is made by turning up the ends and sides. If the volume of the box formed is 756 cu. cm, what are the dimensions of the piece of tin?

pooja195 (pooja195):

The initial dimensions of the piece are a and b, so that a*b = 486 On each side, you cut twice 3 cm, so you cut 6 cm, and the dimensions of the base of the box are a-6 and b-6 The height of the box is of course 3, so that the volume of the box is : (a-6)*(b-6)*3= 756 You have a system of equations ab = 486 3(a-6)(b-6) = 756 Develop the second one (a-6)(b-6) = 756 / 3 = 252 ab -6(a+b) + 36 = 252 ab - 6(a+b) = 252 - 36 = 216 You know that ab = 486, then 486 -6(a+b) = 216 6(a+b) = 486 - 216 = 270 (a+b) = 270 / 6 = 45 Now you have ab = 486 a+b = 45 a = 45 - b (45-b)b = 486 45b - b² = 486 b² -45b + 486 = 0 You solve with the quadratic formula anf you find 18 if b = 18, a = 487/18 = 27 18cm and 27 cm were the dimensions of the tin sheet !

pooja195 (pooja195):

do u get this